Controls Meeting 14-Nov-2013

From GlueXWiki
Jump to: navigation, search


  • Announcements
    • ECS - CSS now connected to cMsg by Graham H
  • Minutes from previous meeting 17-Oct-2013
  • Collimator table wiring - Hovanes, Mark S
  • FCAL installation update - John L
  • Controls installation/testing status - Dave B
  • JInventory update - Serguei P, Elliott
    • BCAL, FCAL, FDC, etc.
  • EPICS update - Hovanes, Vanik
  • Solenoid improvements - Dave, Mark, Scot, Elliott
  • Mantis task tracker
  • Midas elog

  • Future topics
    • Experiment Control Supervisor - Vardan, Elliott


9:30AM Thurs 14-Nov-2013 CC A110


Next Meeting

Note: no meeting Thanksgiving week

9:30am Thurs 12-Dec-2013


Present: Elliott W, Dave B, Beni Z, Hovanes E, Scot S.

PLC Controls

Dave reported on progress on many subsystems:

  • FCAL
    • Controls working, tested as far as possible on the floor, waiting for darkroom completion.
    • Serguei will label everything and handle JInventory database entries.
    • Installation in final location immanent.

  • Gas System
    • PLC mounted on back of gas panel on main platform, much shorter cables needed although need man-lift to work on PLC (uncommon after installation).
    • NBX and Point I/O in gas shed, Ethernet cables will run to switch on FCAL platform.
    • Gas shed plumbing and wiring in progress (by Scot).
    • Currently have two mixing tanks, one each for FDC and CDC.
    • Brooks flow controllers deployed in master/slave configuration.
    • Input to flow controllers 50 psi, 15 psi out.
    • PLC has serial connection to flow controllers.
    • Narcisco modified Beni's original design, followup meeting will be called by Dave B to get final details.

  • BCAL
    • Cables being installed along with four distribution boxes for RTD's.
    • Two CompactLogix PLC's will be used.
    • PLC modules require 3-wire hookup (to 4-wire sensors), must have equal-length cables to maintain precision.
    • Also to be installed is the chiller interface, must be careful about DB-9 to RJ45 adapter wiring.

  • Target
    • Using pulse-tube refrigerator.
    • Lakeshore temperature monitor communication code working (RS232).
    • Lakeshore controller code under development (RS485).
    • One NBX can support one RS485 chain plus one RS232 connection.
    • Control gui will have "Fill" button, no detailed operator procedure needed.
    • Dave Meekins developing fill algorithm which contain sets of setpoints for different conditions.

  • ST
    • Thermocouples read out by target PLC.

  • PS
    • Thermocouples read out by upstream BCAL PLC.

  • Solenoid
    • George maintains list of needed improvements.
    • Dave slowly working on them, e.g: SOE, rewrite of Danfysik communication, etc.


  • Nerses is gone, Vanik continues to work on HV/LV GUIs for upcoming tests.
  • No power to FDC yet, coming soon, need soon!
  • SNMP to MPOS's very slow in hall, ok in EEL, need to figure out what's wrong.
  • May have to use BURT to save/restore setpoints until Acc upgrades MyA to do this.